OpenStudy (whatever):

an airplane flew for 5 hours with a tail wind of 25 mi/h. The return flight against the same wind took 5.5 hours. ind the speed of the airplane in air.

OpenStudy (dumbcow):

d = r*t The trip is the same distance there and back, so set distances equal to each other. Speed is increased by having tail wind and decreased by going against the wind. Let r be speed of airplane. 5(r+25) = 5.5(r-25) 5r + 125 = 5.5r - 137.5 -0.5r = -262.5 r = 525